# Near-Optimal Decision Trees in a SPLIT Second

### v2

@@ A.7.2 Preprocessing
  • All datasets have a combination of categorical and continuous features, while SPLIT / LicketySPLIT / RESPLIT and many other decision tree algorithms require binarization of features. We therefore use the threshold guessing mechanism of binarization from McTavish et al. (2022), which can handle both these feature types. In particular:
@@
    • For version $1$, we chose $n_{est}$ for each dataset such that the resulting binarized dataset has between $40$-$100$ features. This is the version used for experiments in Figures 4, 5, 6, 7 when we compare SPLIT / LicketySPLIT with other datasets.
    • For version $2$, we chose $n_{est}$ for each dataset such that the resulting dataset has around $20$-$25$ features. This is the version used when we use the TreeFARMS algorithm (Xin et al., 2022) to generate Rashomon sets to explore the properties of near optimal decision trees, as TreeFARMS can be very slow otherwise. Figures 2 and 8 use this version of the datasets.
@@
Data Set Samples ##\## Features ##\## Features After Binarization
##\## Features After
@@
Table 3: Characteristics of the $9$ datasets tested in this paper for LicketySPLIT and SPLIT experiments. We generate two binarized versions of each dataset using the threshold guessing mechanism in (McTavish et al., 2022) which are used for different sets of experiments.
Data Set Samples ##\## Features ##\## Features After Binarization
@@ A.7.3 Details of Comparative Experiments for SPLIT and LicketySPLIT
  • Greedy: This is the standard scikit-learn DecisionTreeClassifier class that implements CART. We vary the sparsity of this algorithm by changing the min__\__samples__\__leaves argument. This is the minimum number of examples required to be in a leaf in order for CART to make further a split at that point.
  • GOSDT (Lin et al., 2020): We vary the sparsity parameter $\lambda$, choosing equispaced values from $0.001$ to $0.02$.
@@
  • Thompson Sampled Decision Trees (TSDT) (Chaouki et al., 2024): Following the practices described in the Appendix Section $B$ of their paper, we fix the following parameters:
@@
    • thresh__\__tree $=-1e-6$
    • thresh__\__leaf $=1e-6$
    • thresh__\__mu $=0.8$
    • thresh__\__sigma $=0.1$
@@
  • Murtree (Demirović et al., 2022): For this method, we vary the max__\__num__\__nodes, which is the hard sparsity constraint imposed by Murtree on the number of leaves.
    • For depth $5$ trees, we choose max__\__num__\__nodes in the set $\{4,5,6,7,8,9,10,11\}$.
    • For depth $4$ trees, we choose max__\__num__\__nodes in the set $\{3,4,5,6,7\}$.
  • MAPTree (Sullivan et al., 2024): The paper has two hyperparameters, $\alpha$ and $\beta$, which in theory control for sparsity in theory by adjusting the prior. However, the authors show that MAPTree does not exhibit significant sensitivity to $\alpha$ and $\beta$ across any metric. Therefore, the only parameter we choose to vary for this experiment is num__\__expansions. We chose $10$ values of this parameter in a logarithmically spaced interval from $[10^{0},10^{3.5}]$.
  • Top-k (DL8.5) (Blanc et al., 2024): The paper also does not specify how to vary the sparsity parameter - hence, we vary $k$ from $1$-$10$.
@@
We briefly discuss another decision tree optimization algorithm from (Balcan & Sharma, 2024) that demonstrates good performance on a tabular dataset. This method first proposes a novel greediness criterion called the $(\alpha,\beta)$-Tsallis entropy, defined as:
@@
For the COMPAS dataset, which is one of the smaller datasets in our experiments, we conducted a brief evaluation by averaging results over $3$ trials for $3$ different values of the hyperparameters $\alpha$ and $\beta$, arranged in a grid-based configuration as defined in (Balcan & Sharma, 2024). These hyperparameters influence the functional form of the above greedy heuristic. Below, we summarize the key settings and observations:
@@
    The method in (Balcan & Sharma, 2024) achieves approximately 31.6% test error with around 10 leaves, requiring an average of 10 minutes to train for a single hyperparameter setting. Another thing to note is that it isn’t clear a-priori which hyperparameter will lead to the best performance (in terms of the desired objective in Equation 1), so many different combinations of hyper-parameters might need to be tested in order to find a well-performing tree.
@@
In summary, our proposed methods are over 600x faster than (Balcan & Sharma, 2024), with a negligible difference in test performance.
A Note on Comparative Experiments with Blossom (Demirović et al., 2023)
@@
Table 5: Comparison of Blossom and LicketySPLIT across datasets when Blossom is allowed to finish. We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across $5$ trials with depth budget $5$). Only the mean is bolded when LicketySPLIT performs better. Values are reported as mean [lower, upper] (indicating 95% confidence intervals).
From Table 5, we see that, despite being allowed to run to completion (taking over 10×10$\times$10 × longer in many cases), Blossom often underperforms LicketySPLIT in test loss. Furthermore, it is much less sparse than LicketySPLIT, having over 4×4$\times$4 × as many leaves for similar test performances.
@@
Table 6: Comparison of Blossom and LicketySPLIT in an anytime setting (i.e. Blossom execution is stopped around the same time as LicketySPLIT) . We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across $5$ trials with depth budget $5$). Only the mean is bolded where LicketySPLIT outperforms Blossom. Values are reported as mean [lower, upper] (indicating $95\%$ confidence intervals). Note that the Blossom algorithm is not able to explicitly account for sparsity, hence it always returns fully grown trees up to a given depth.
@@ A.7.4 Description of Machines Used
All experiments were performed on an institutional computing cluster. This was a single Intel(R) Xeon(R) Gold 6226 machine with a $2.70$GHz CPU. It has $300$GB RAM and $24$ cores.
@@ A.8 Appendix Proofs
Figure 18: The Lambert W function, which has two branches in the real plane, $W_{0}(z)$ and $W_{-1}(z)$. Figure from (Lóczi, 2021).
@@
(Chatzigeorgiou, 2013) show the following lower bound for $W_{-1}(z)$:
@@
(Lóczi, 2021) show the following upper bound for $W_{-1}(z)$:
@@
$=\frac{\log\left(2e^{2}k^{-\left(d+1\right)}\log k\right)-\log\Big(-\log\left(-2e^{2}k^{-\left(d+1\right)}\log k\right)\Big{missing})-2+2d\log k}{2\log k}$ (45)
$=\frac{\log 2-(d+1)\log k+\log\log k-\log\Big(-\big{(}\log 2+2-\left(d+1\right)\log k+\log\log k\big{)}\Big{missing})+2d\log k}{2\log k}.$ (46)
@@
$\log\Big(\left(d+1\right)\log k-\log\log k\Big{missing})\geq 1-\frac{1}{\left(d+1\right)\log k-\log\log k}$ (48)
@@
This is clearly $>0$ as the $\log^{2}k$ terms dominate $\log k$. Thus, the value $d_{l}^{*}=\frac{d-1}{2}$ corresponds to the minimum of the runtime.
@@
$\log\Gamma(d_{l}+1)-\log\Gamma(d-d_{l}+1)-\log(d-d_{l})=\log\Big(\frac{d_{l}!}{(d-d_{l})!(d-d_{l})}\Big{missing}).$ (78)
Notice that for any $d_{l}$ between 00 and $\lfloor\frac{d}{2}\rfloor$ (inclusive), the RHS is always less than 00. Similarly, for $\lfloor\frac{d}{2}\rfloor<d_{l}\leq d-1$, the term is always greater than 00. Given that these are constant as $k$ increases:
@@
$-\mathcal{O}\Big{(}\frac{1}{\log k}\Big{)}\leq\frac{\log\Gamma(d_{l}+1)-\log\Gamma(d-d_{l}+1)-\log(d-d_{l})}{\log k}\leq\mathcal{O}\Big{(}\frac{1}{\log k}\Big{)}$ (80)
for all $10\leq d_{l}\leq d-$ 1 (which are the constraints in our setup). Hence, we conclude that, for large $k$:
@@
We show that the computed value of $d_{l}^{*}$ is indeed the minimum for large $k$ by evaluating the second derivative of the runtime, i.e. $\frac{\partial^{2}}{\partial d_{l}^{2}}\Big{(}\frac{n(d-d_{l})k^{d_{l}+1}}{d_{l}!}+\frac{nk^{d-d_{l}}}{(d-d_{l})!}\Big{)}\evaluated{}_{d_{l}=d_{l}^{*}}$.
@@
$=\Bigg{[}\frac{\partial}{\partial d_{l}}\Big{(}\log k(d-d_{l})k^{d_{l}+1}\Big{)}\cdot\frac{\Gamma(d_{l}+1)}{\Gamma(d_{l}+1)^{2}}-\frac{\log k(d-d_{l})k^{d_{l}+1}}{\Gamma(d_{l}+1)^{2}}\cdot\frac{\partial}{\partial d_{l}}\Gamma(d_{l}+1)\Bigg{]}\ \ -$ (87)
@@
$=\Bigg{[}\Big{(}-\log kk^{d_{l}+1}+(d-d_{l})k^{d_{l}+1}(\log k)^{2}\Big{)}\cdot\frac{1}{\Gamma(d_{l}+1)}-\frac{\log k(d-d_{l})k^{d_{l}+1}}{\Gamma(d_{l}+1)^{2}}\cdot\Gamma(d_{l}+1)\psi(d_{l}+1)\Bigg{]}\ \ -$ (89)
@@
$=\Bigg{[}\frac{\Big{(}-\log kk^{d_{l}+1}+(d-d_{l})k^{d_{l}+1}(\log k)^{2}\Big{)}}{\Gamma(d_{l}+1)}-\frac{\log k(d-d_{l})k^{d_{l}+1}\psi(d_{l}+1)}{\Gamma(d_{l}+1)}\Bigg{]}\ \ -$ (91)
@@
$=\Bigg{[}\frac{\Big{(}-k^{d_{l}+1}(\log k)+(d-d_{l})k^{d_{l}+1}(\log k)^{2}-\log k(d-d_{l})k^{d_{l}+1}\psi(d_{l}+1)\Big{)}}{\Gamma(d_{l}+1)}\Bigg{]}\ \ -$ (93)
@@
[…]
@@ A.8.5 Additional Claims, with Proofs
Our proof follows a similar construction as (Blanc et al., 2024). They define the function Tribes as follows:
Definition A.4 (Tribes: from Blanc et al. 2024).
@@
where $t=\Big{\lfloor}\frac{\ell}{w}\Big{\rfloor}$. Blanc et al. (2019) prove the following properties of Tribes:
@@
[…]
@@
We now apply Lemma $4.4$ from (Blanc et al., 2024) in this context, adjusting the notation to suit our case. Let $T$ be the tree of depth $d$ returned by greedy. Consider any root-to-leaf path of $T$ that does not query any of the first $d_{l}$ features of $\mathbf{x}$ (i.e. the Tribes block). Only features from the Majority block are therefore queried by $T$ along this path. We can therefore write the probability of error along this path:
@@
where the last line follows, because Tribes is nearly balanced. As the distribution over x is uniform, each leaf is equally likely. (Blanc et al., 2024) then show that, if only $p$-fraction of root-to-leaf paths of $T$ query at least one of the first $d_{l}$ coordinates, then:
@@
We now want to show that, just like in the case of (Blanc et al., 2024), $p$ is small asymptotically. If this is the case, we can claim that a greedy tree is arbitrarily bad. The only difference between (Blanc et al., 2024) and us is that their greedy tree has depth $d_{l}$ (adjusting for notation), but we want to construct a tree of depth $d$.
We now use Lemma $7.4$ from (Blanc et al., 2019), which proves the following (again, adjusting for our notation): A random root-to-leaf path of a greedy tree $T$ satisfies the following with probability at least $1-\mathcal{O}(u^{-2})$: If the length of this path is less than $\mathcal{O}(\frac{u}{\log u})$, at any point along that path, all coordinates within the majority block that have not already been queried have correlation at least $\frac{1}{100\sqrt{u}}$. Now, for a greedy tree of depth $d$:
@@
  • Sub-block $1$ corresponds to the $d_{l}$ features for which we will compute a parity bit. At a high level, a tree needs to know the parity of the expression in order to ‘unlock’ a high accuracy. This also serves to ‘trick’ greedy into not choosing these features, because they will have 00 correlation with the label. Let $X_{1}\ldots X_{d_{l}}$ be the features in this sub-block.
@@
We can immediately see that the best tree will achieve an error $\geq 1-\epsilon$. The Rashomon set in this case will contain $R-1$ trees (besides the empirical risk minimizer). In particular, each tree $T$ in the Rashomon set will split on one unique feature $X_{j}\ \forall 2d<j\leq 2d+R$, making a prediction on an instance $\mathbf{X}=(X_{1},...X_{k})$ of the following form:
@@
A greedy tree will seek to split on the feature that has the highest correlation with the label $y$. From the definition of $y$ in Equation 112, it follows that $X_{d+1},\ldots,X_{2d}$ are the only variables that will have non 00 correlation with the label outcome. Thus, the tree will fully split only on these features up to depth $d$. However, this means that the tree does not learn the underlying parity function $(X_{1}\oplus\ldots\oplus X_{d_{l}})$. Thus, $1-\epsilon$ proportion of the time, the tree will achieve $\frac{1}{2}$ accuracy. Thus, the total accuracy is less than $\frac{1}{2}(1-\epsilon)+\epsilon=\frac{1}{2}+\epsilon$. ∎
@@
$y=\begin{cases}x_{1}\oplus\textrm{Majority}(x_{2},\ldots x_{d})&\textrm{with probability $1-\epsilon$}\\ \textrm{Majority}(x_{d+1},\ldots x_{2d})&\textrm{with probability $\epsilon$}\\ \end{cases}$

### v3

@@ A.7.2 Preprocessing
  • All datasets have a combination of categorical and continuous features, while SPLIT / LicketySPLIT / RESPLIT and many other decision tree algorithms require binarization of features. We therefore use the threshold guessing mechanism of binarization from gosdt_guesses, which can handle both these feature types. In particular:
@@
    • For version $1$, we chose $n_{est}$ for each dataset such that the resulting binarized dataset has between 4040-$100$ features. This is the version used for experiments in Figures 4, 5, 6, 7 when we compare SPLIT / LicketySPLIT with other datasets.
    • For version $2$, we chose $n_{est}$ for each dataset such that the resulting dataset has around 2020-$25$ features. This is the version used when we use the TreeFARMS algorithm (xin2022treefarms) to generate Rashomon sets to explore the properties of near optimal decision trees, as TreeFARMS can be very slow otherwise. Figures 2 and 8 use this version of the datasets.
@@
Data Set Samples #\# Features #\# Features After Binarization
#\# Features After
@@
Table 3: Characteristics of the $9$ datasets tested in this paper for LicketySPLIT and SPLIT experiments. We generate two binarized versions of each dataset using the threshold guessing mechanism in (gosdt_guesses) which are used for different sets of experiments.
Data Set Samples #\# Features #\# Features After Binarization
@@ A.7.3 Details of Comparative Experiments for SPLIT and LicketySPLIT
  • Greedy: This is the standard scikit-learn DecisionTreeClassifier class that implements CART. We vary the sparsity of this algorithm by changing the min_\_samples_\_leaves argument. This is the minimum number of examples required to be in a leaf in order for CART to make further a split at that point.
  • GOSDT (gosdt): We vary the sparsity parameter $\lambda$, choosing equispaced values from $0.001$ to $0.02$.
@@
  • Thompson Sampled Decision Trees (TSDT) (thompson_aistats): Following the practices described in the Appendix Section $B$ of their paper, we fix the following parameters:
@@
    • thresh_\_tree =1e6=-1e-6
    • thresh_\_leaf =1e6=1e-6
    • thresh_\_mu =0.8=0.8
    • thresh_\_sigma =0.1=0.1
@@
  • Murtree (murtree): For this method, we vary the max_\_num_\_nodes, which is the hard sparsity constraint imposed by Murtree on the number of leaves.
    • For depth $5$ trees, we choose max_\_num_\_nodes in the set $\{4,5,6,7,8,9,10,11\}$.
    • For depth $4$ trees, we choose max_\_num_\_nodes in the set $\{3,4,5,6,7\}$.
  • MAPTree (maptree): The paper has two hyperparameters, $\alpha$ and $\beta$, which in theory control for sparsity in theory by adjusting the prior. However, the authors show that MAPTree does not exhibit significant sensitivity to $\alpha$ and $\beta$ across any metric. Therefore, the only parameter we choose to vary for this experiment is num_\_expansions. We chose $10$ values of this parameter in a logarithmically spaced interval from $10^{0},10^{3.5}]$.
  • Top-k (DL8.5) (topk): The paper also does not specify how to vary the sparsity parameter - hence, we vary $k$ from 11-$10$.
@@
We briefly discuss another decision tree optimization algorithm from (slow_greedy) that demonstrates good performance on a tabular dataset. This method first proposes a novel greediness criterion called the $(\alpha,\beta)$-Tsallis entropy, defined as:
@@
For the COMPAS dataset, which is one of the smaller datasets in our experiments, we conducted a brief evaluation by averaging results over $3$ trials for $3$ different values of the hyperparameters $\alpha$ and $\beta$, arranged in a grid-based configuration as defined in (slow_greedy). These hyperparameters influence the functional form of the above greedy heuristic. Below, we summarize the key settings and observations:
@@
    The method in (slow_greedy) achieves approximately 31.6% test error with around 10 leaves, requiring an average of 10 minutes to train for a single hyperparameter setting. Another thing to note is that it isn’t clear a-priori which hyperparameter will lead to the best performance (in terms of the desired objective in Equation 1), so many different combinations of hyper-parameters might need to be tested in order to find a well-performing tree.
@@
In summary, our proposed methods are over 600x faster than (slow_greedy), with a negligible difference in test performance.
A Note on Comparative Experiments with Blossom (demirovic2023blossom)
@@
Table 5: Comparison of Blossom and LicketySPLIT across datasets when Blossom is allowed to finish. We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across 55 trials with depth budget 55). Only the mean is bolded when LicketySPLIT performs better. Values are reported as mean [lower, upper] (indicating 95% confidence intervals).
From Table 5, we see that, despite being allowed to run to completion (taking over $10\times$ longer in many cases), Blossom often underperforms LicketySPLIT in test loss. Furthermore, it is much less sparse than LicketySPLIT, having over $4\times$ as many leaves for similar test performances.
@@
Table 6: Comparison of Blossom and LicketySPLIT in an anytime setting (i.e. Blossom execution is stopped around the same time as LicketySPLIT) . We show the LicketySPLIT configuration (after grid-search across $\lambda$) that yielded the best test loss on that dataset (averaged across 55 trials with depth budget 55). Only the mean is bolded where LicketySPLIT outperforms Blossom. Values are reported as mean [lower, upper] (indicating $95\%$ confidence intervals). Note that the Blossom algorithm is not able to explicitly account for sparsity, hence it always returns fully grown trees up to a given depth.
@@ A.7.4 Description of Machines Used
All experiments were performed on an institutional computing cluster. This was a single Intel(R) Xeon(R) Gold 6226 machine with a 2.702.70GHz CPU. It has 300300GB RAM and $24$ cores.
@@ A.8 Appendix Proofs
Figure 18: The Lambert W function, which has two branches in the real plane, $W_{0}(z)$ and $W_{-1}(z)$. Figure from (lambert).
@@
(lambert_bound) show the following lower bound for $W_{-1}(z)$:
@@
(lambert) show the following upper bound for $W_{-1}(z)$:
@@
$=\frac{\log\left(2e^{2}k^{-\left(d+1\right)}\log k\right)-\log\Big(-\log\left(-2e^{2}k^{-\left(d+1\right)}\log k\right)\Big.)-2+2d\log k}{2\log k}$ (45)
$=\frac{\log 2-(d+1)\log k+\log\log k-\log\Big(-\big(\log 2+2-\left(d+1\right)\log k+\log\log k\big)\Big.)+2d\log k}{2\log k}.$ (46)
@@
$\log\Big(\left(d+1\right)\log k-\log\log k\Big.)\geq 1-\frac{1}{\left(d+1\right)\log k-\log\log k}$ (48)
@@
This is clearly >0>0 as the $\log^{2}k$ terms dominate $\log k$. Thus, the value $d_{l}^{*}=\frac{d-1}{2}$ corresponds to the minimum of the runtime.
@@
[…]
Notice that for any $d_{l}$ between 0 and $\lfloor\frac{d}{2}\rfloor$ (inclusive), the RHS is always less than 0. Similarly, for $\lfloor\frac{d}{2}\rfloor<d_{l}\leq d-1$, the term is always greater than 0. Given that these are constant as $k$ increases:
@@
$\mathcal{O}\Big(\frac{1}{\log k}\Big)\leq\frac{\log\Gamma(d_{l}+1)-\log\Gamma(d-d_{l}+1)-\log(d-d_{l})}{\log k}\leq\mathcal{O}\Big(\frac{1}{\log k}\Big)$ (80)
for all $0\leq d_{l}\leq d-1$ (which are the constraints in our setup). Hence, we conclude that, for large $k$:
@@
[…]
@@
$=\Bigg[\frac{\partial}{\partial d_{l}}\Big(\log k(d-d_{l})k^{d_{l}+1}\Big)\cdot\frac{\Gamma(d_{l}+1)}{\Gamma(d_{l}+1)^{2}}-\frac{\log k(d-d_{l})k^{d_{l}+1}}{\Gamma(d_{l}+1)^{2}}\cdot\frac{\partial}{\partial d_{l}}\Gamma(d_{l}+1)\Bigg]\ \$- (87)
@@
$=\Bigg[\Big(-\log kk^{d_{l}+1}+(d-d_{l})k^{d_{l}+1}(\log k)^{2}\Big)\cdot\frac{1}{\Gamma(d_{l}+1)}-\frac{\log k(d-d_{l})k^{d_{l}+1}}{\Gamma(d_{l}+1)^{2}}\cdot\Gamma(d_{l}+1)\psi(d_{l}+1)\Bigg]\ \$- (89)
@@
$=\Bigg[\frac{\Big(-\log kk^{d_{l}+1}+(d-d_{l})k^{d_{l}+1}(\log k)^{2}\Big)}{\Gamma(d_{l}+1)}-\frac{\log k(d-d_{l})k^{d_{l}+1}\psi(d_{l}+1)}{\Gamma(d_{l}+1)}\Bigg]\ \$- (91)
@@
$=\Bigg[\frac{\Big(-k^{d_{l}+1}(\log k)+(d-d_{l})k^{d_{l}+1}(\log k)^{2}-\log k(d-d_{l})k^{d_{l}+1}\psi(d_{l}+1)\Big)}{\Gamma(d_{l}+1)}\Bigg]\ \$- (93)
@@
Note that the $(\log k)^{2}$ terms are dominant in this expression as k $\rightarrow\infty$. Hence, at $d_{l}^{*}=\frac{d-1}{2}$, the terms that will affect the sign of the expression are:
@@ A.8.5 Additional Claims, with Proofs
Our proof follows a similar construction as (topk). They define the function Tribes as follows:
Definition A.4 (Tribes: from topk).
@@
where $t=\Big\lfloor\frac{\ell}{w}\Big\rfloor$. blanc2019top prove the following properties of Tribes:
@@
$\text{Maj}(x):=\mathbf{1}[\text{at least half of }x\text{'s coordinates are }1].$
@@
We now apply Lemma $4.4$ from (topk) in this context, adjusting the notation to suit our case. Let $T$ be the tree of depth $d$ returned by greedy. Consider any root-to-leaf path of $T$ that does not query any of the first $d_{l}$ features of $\mathbf{x}$ (i.e. the Tribes block). Only features from the Majority block are therefore queried by $T$ along this path. We can therefore write the probability of error along this path:
@@
where the last line follows, because Tribes is nearly balanced. As the distribution over x is uniform, each leaf is equally likely. (topk) then show that, if only $p$-fraction of root-to-leaf paths of $T$ query at least one of the first $d_{l}$ coordinates, then:
@@
We now want to show that, just like in the case of (topk), $p$ is small asymptotically. If this is the case, we can claim that a greedy tree is arbitrarily bad. The only difference between (topk) and us is that their greedy tree has depth $d_{l}$ (adjusting for notation), but we want to construct a tree of depth $d$.
We now use Lemma $7.4$ from (blanc2019top), which proves the following (again, adjusting for our notation): A random root-to-leaf path of a greedy tree $T$ satisfies the following with probability at least $1-\mathcal{O}(u^{-2})$: If the length of this path is less than $\mathcal{O}(\frac{u}{\log u})$, at any point along that path, all coordinates within the majority block that have not already been queried have correlation at least $\frac{1}{100\sqrt{u}}$. Now, for a greedy tree of depth $d$:
@@
  • Sub-block $1$ corresponds to the $d_{l}$ features for which we will compute a parity bit. At a high level, a tree needs to know the parity of the expression in order to ‘unlock’ a high accuracy. This also serves to ‘trick’ greedy into not choosing these features, because they will have 0 correlation with the label. Let $X_{1}\ldots X_{d_{l}}$ be the features in this sub-block.
@@
We can immediately see that the best tree will achieve an error $\geq 1-\epsilon$. The Rashomon set in this case will contain R1R-1 trees (besides the empirical risk minimizer). In particular, each tree $T$ in the Rashomon set will split on one unique feature $X_{j}\ \forall 2d<j\leq 2d+R$, making a prediction on an instance $\mathbf{X}=(X_{1},...X_{k})$ of the following form:
@@
A greedy tree will seek to split on the feature that has the highest correlation with the label $y$. From the definition of $y$ in Equation 112, it follows that $X_{d+1},\ldots,X_{2d}$ are the only variables that will have non 0 correlation with the label outcome. Thus, the tree will fully split only on these features up to depth $d$. However, this means that the tree does not learn the underlying parity function $(X_{1}\oplus\ldots\oplus X_{d_{l}})$. Thus, $1-\epsilon$ proportion of the time, the tree will achieve $\frac{1}{2}$ accuracy. Thus, the total accuracy is less than $\frac{1}{2}(1-\epsilon)+\epsilon=\frac{1}{2}+\epsilon$. ∎
@@
[…]